OpenStudy (anonymous):

how many different ways can 2 balls be drawn from a bag containing 5 balls? each ball is a different color, and first ball is replaced before the second one is selected. also how many ways are there if the ball is not replaced?

OpenStudy (kropot72):

The number of permutations of n different things taken k at a time with repetitions is: \[n ^{k}=5^{2}=you\ can\ calculate\] The number of permutations of n different things taken k at a time without repetitions is: \[\frac{n!}{(n-k)!}=\frac{5!}{(5-2)!}=you\ can\ calculate\]

OpenStudy (ybarrap):

Case 1 A. How many ways can you pick the 1st ball: 5 ways B. Put the ball back in the bag C. How many ways can you pick the 2nd ball: 5 ways The number of ways of doing A and C is 5*5=25 Case 2 A. How many ways can you pick the 1st ball: 5 ways B. DON'T put the ball back in the bag C. How many ways can you pick the 2nd ball: 4 ways The number of ways of doing A and C is 5*4=20 Once you understand the concept, you can move to formulas to make the calculations easier.